OpenStudy (anonymous):

The population of a certain island has an annual growth rate of 6.9%. If you start with a poplulation of 2500 people, what would the population be in 10 years? Round your answer to the nearest hundred. The population would be about

OpenStudy (anonymous):

@study100

OpenStudy (anonymous):

4,872 is incorrect

OpenStudy (anonymous):

@Power2Knowledge Sorry I wonder if you're learnning exponential growth or compound anual growth? I think this is exponential growth: y= Ce^(rt) y= 2500 e^(0.069*10) y=4984 people

OpenStudy (anonymous):

exponential growth

OpenStudy (anonymous):

so there it is :)

OpenStudy (anonymous):

thank you. can you help me with a few more questions?

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

What are the solutions to the equation https://wca.sooschools.com/media/g_alg02_ccss_2013/9/rad5xpls1.gif = x + 1? https://wca.sooschools.com/media/g_alg02_ccss_2013/9/img_alg2_9_5a_04_graphing.gif The solutions are x =

OpenStudy (anonymous):

(i received partial credit) but my answer is 5,0

OpenStudy (anonymous):

is this the whole equation for number 1? \[\sqrt{5x+1}= x+1\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for number 2, x= where they touch so x=0,3

OpenStudy (anonymous):

square both sides for number 1 5x+1 = x^2 + 2x +1 x^2 -3x= 0 x(x-3)=0 x= 0,3 same answer

OpenStudy (anonymous):

Thanks. What are the solutions to the equation 4x^3 - 5x = |4x|? List your answers in increasing order. https://wca.sooschools.com/media/g_alg02_ccss_2013/9/img_alg2_9_5a_06_graphing.gif The solutions are x = my answer (partial) 3/2, -3/2, 0